Solve the equation..
-4/21y = -8/15

Answer by stanbon(75887)

-4/21y = -8/15
----
y = (-21/4)(-8/15)
----
y = (7/1)(2/5)
---
y = 14/5
---
y = 2.8
============
Cheers,
Stan H.